Self- and Foreign Gas Broadening in the ν_{1}+ν_{3} Band of SO_{2}*

\underline{B. SUMPF**}, D. REHLE, H.-D. KRONFELDT, Optisches Institut der Technischen Universit t Berlin, Sekr. PN 0-1, Hardenbergstr. 36, 10623 Berlin, Germany; V.V. LAZAREV***, Institute of Atmospheric Optics, Siberian Branch, Russian Academy of Sciences, 1 Academicheskii av., Tomsk 634055 Russia.

Experimental results concerning the self-broadening of about 100 lines in the ν_{1}+ν_{3} band of SO_{2} near 2499 cm^{−1} will be presented. The experiments were carried out with a tunable diode laser spectrometer in pulsed mode. The dependence of the broadening coefficient on the rotational quantum number J′′ and K′′_{a} will be presented together with the comparison to our former results concerning the ν_{1} and ν_{3} band, respectively. In the studied quantum number range 2 ≤ J′′ ≤ 47, 0 ≤ Ka′′ ≤ 13 we observe similar to the other bands a pronounced K′′_{a} dependence but only a weak J′′ dependence. The investigations are completed by measurements with the perturbers He, Ne, Ar, Kr, Xe, N_{2}, O_{2}, H_{2}, and D_{2} for some selected lines.

* This project was supported by the Deutsche Forschungsgemeinschaft
** B.S. acknowledges a grant from the Deutsche Forschungsgemeinschaft
*** V.V.L. acknowledges a grant from the Deutscher Akademischer Austauschdienst